OpenStudy (anonymous):

Which statement is true for the equation 3x − 3x − 2 = −2?

OpenStudy (anonymous):

where are the statements?

OpenStudy (anonymous):

It has no solution. It has one solution. It has two solutions. It has infinitely many solutions.

OpenStudy (anonymous):

hello.........

Nnesha (nnesha):

try to solve it first if you get equal sides then infinitely many solution example 4=4 if both sides are not equal then no solution like 9=7

Nnesha (nnesha):

if you get x = something then one solution

Nnesha (nnesha):

\[\huge\rm \color{blue}{3x − 3x }\color{reD}{− 2 = −2}\]combine like terms

OpenStudy (decentnabeel):

3x − 3x − 2 = −2 no solution add 2 on both sides 3x-3x=-2+2 0=0

Nnesha (nnesha):

nice try ...

Nnesha (nnesha):

but guess what

OpenStudy (anonymous):

what

Nnesha (nnesha):

what's your answer ?

OpenStudy (anonymous):

no solution

Nnesha (nnesha):

alright then.

OpenStudy (anonymous):

is it right

Nnesha (nnesha):

I believe you didn't read my first comment

OpenStudy (anonymous):

wdy mean

OpenStudy (anonymous):

@sweetburger can you check this out

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha if you get equal sides then infinitely many solution example 4=4 if both sides are not equal then no solution like 9=7 \(\color{blue}{\text{End of Quote}}\)

Nnesha (nnesha):

3x-3x = ?

OpenStudy (anonymous):

0

Nnesha (nnesha):

so you get equal sides \[\huge\rm \color{blue}{\cancel{3x − 3x} }\color{reD}{− 2 = −2}\] -2=-2 so answer is ?

OpenStudy (anonymous):

o

Nnesha (nnesha):

what's that ?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha if you get equal sides then infinitely many solution example 4=4 if both sides are not equal then no solution like 9=7 \(\color{blue}{\text{End of Quote}}\) you got -2=-2 so which one is correct answer?

OpenStudy (anonymous):

one solution

OpenStudy (decentnabeel):

so @Nnesha what is the value of x.. you say many solution

OpenStudy (anonymous):

i am confused

Nnesha (nnesha):

if you get equal sides then infinitely many solutions if both sides are not equal then no solution

Nnesha (nnesha):

you got -2=-2 both sides are equal so answer is ?

OpenStudy (decentnabeel):

only tell me what is the value of x..

OpenStudy (anonymous):

so there is one solution

Nnesha (nnesha):

no

OpenStudy (anonymous):

it is zero

Nnesha (nnesha):

x cancels out right you get -2=-2 equal sides

Nnesha (nnesha):

let me repeat t if both sides are equal like 4=4 then your answer would be infinitely many solutions

OpenStudy (decentnabeel):

@Nnesha you mean no value of x.. you said previous comment in many solution.. so what solution

OpenStudy (anonymous):

oh

Nnesha (nnesha):

4=9 no solution x= something , one solution 4=4 infinitely many solution

Nnesha (nnesha):

so what do you think ?

OpenStudy (anonymous):

many solutions

Nnesha (nnesha):

thats right!

OpenStudy (anonymous):

thank you

Nnesha (nnesha):

Yw

OpenStudy (decentnabeel):

no the answer is.. no solution

Nnesha (nnesha):

how is it no solution ?

OpenStudy (decentnabeel):

3x-3x=0 so that the variable is the cancle out

Nnesha (nnesha):

okay you will get 0=0 so thats mean both sides are equal

OpenStudy (anonymous):

soooo

Nnesha (nnesha):

so what ?

OpenStudy (decentnabeel):

then solve it.. so what is the variable for solving.. there is no variable for sloving.. so there is no solution

Nnesha (nnesha):

......

Nnesha (nnesha):

repeating when you get equal sides 't the end then infinitely many solution believe it or NOT

OpenStudy (decentnabeel):

not

Nnesha (nnesha):

as you wish, sir

OpenStudy (decentnabeel):

AS you wsih @Nnesha expert

Nnesha (nnesha):

you can tag some expert here i wouldn't mind ..

OpenStudy (decentnabeel):

yes i have no mind.. i clear the concept

OpenStudy (decentnabeel):

This solution could also be stated as "all real numbers" or "all reals" or "the whole number line"; expect some variation in lingo from one text to the next. Note that, if I had solved the equation by subtracting a 5 from either side of 5 + 4x = 5 + 4x to get "4x = 4x", I would have ended up with nothing other than another trivially-true statement. I could also have subtracted both 4x and 5 from both sides to get "0 = 0", but the solution would still be the same: "all x". Don't be surprised if, for "all real numbers" or "no solution" equations, you don't necessarily have the exact same steps as some of your fellow students. Since there are infinitely-many always-true equations (like "0 = 0") and infinitely-many nonsensical equations (like "3 = 4"), there will be many ways of arriving at these answers

Nnesha (nnesha):

you can substitute x by any number you will get equal sides \[\huge\rm 3(9)-3(9)-2=-2\] or let plug in 2 \[\huge\rm 3(-2)-3(-2)-2=-2\] both sides are e qual how can you say that there is no solution ?

OpenStudy (decentnabeel):

i think both are true.. some expert say no solution.. and some one is many

Nnesha (nnesha):

well you can plug in any number you will get equal sides so thats means infinitely many solution

OpenStudy (decentnabeel):

http://www.purplemath.com/modules/solvelin4.htm @Nnesha expert you can see it

Nnesha (nnesha):

so http://prntscr.com/7yn3yb solution is all x thats mean any number can be x

Nnesha (nnesha):

http://prntscr.com/7yn46i

OpenStudy (decentnabeel):

so the finally answer is.. this is no solution for this equation.. but many solution for all x

OpenStudy (decentnabeel):

@Nnesha are you clear

Nnesha (nnesha):

wait what ?

Nnesha (nnesha):

x can be any number so infinitely many solution :=)

OpenStudy (decentnabeel):

yes i agree many solution for x.. but there is no solution for equation

Nnesha (nnesha):

what do you mean ? i already posted some example where we get equal sides so thats mean there are many solution for equation

OpenStudy (decentnabeel):

ohhhhhhhhhhhhhhhh

OpenStudy (decentnabeel):

as you wish

Nnesha (nnesha):

as you wish ? no thats how it works *mathematically*